OpenStudy (anonymous):

prove sec(x)-1 / sec(x)+1 + cos(x)-1/cos(x)+1

jimthompson5910 (jim_thompson5910):

i think you meant to say \[\Large \frac{\sec(x)-1}{\sec(x)+1} = \frac{\cos(x)-1}{\cos(x)+1}\] right?

OpenStudy (xapproachesinfinity):

you have + or = btw the two fractions?